# Forecasting the future development in quality and value of professional football players for applications in team management

### v1

@@ 2.2 Predicting future values
Although many studies in sports analytics are concerned with the quantification of the quality and monetary value of football players at the present moment, only limited work has been done on the future development of player performance. Apostolou and Tjortjis [34] have conducted a small-scale study with the aim to predict the number of future goals of the two football players Lionel Messi and Luis Suárez using a random forest, logistic regression, a multi-layer perception classifier, and a linear support vector classifier. Pantzalis and Tjortjis [35] have predicted the expert ratings in the next season of 59 center-backs in the English Premier League based on one season of player attributes from a popular football manager simulation game. Their method uses a linear regression model to describe the in-sample patterns. Giannakoulas et al. [36] have trained linear regression, random forest, and multi-layer perceptron models to predict the number of goals of a football playe rin a season before the start of the corresponding season. Their dataset entails around 800 football players. Similarly, the models by Markopoulou et al. [37] have been trained to predict the number of goals by looking at the creation of different models per competition. This study on 424 football players shows that the best results are often obtained using XGBoost models for this prediction problem and that making different models for different competitions might be beneficial.
@@
Our research treats forecasting of both player quality and transfer value. The current paper builds on the existing literature about forecast player performance by studying the long-term forecasting of a model-based player quality indicator on a larger data set. Additionally, this research avoids the bias in current literature that better describes offensive players by using a more general top-down rating. Our research contributes to the existing knowledge of forecasting the monetary value because the models are trained to perform out-of-sample prediction, which makes it possible to apply it to unseen situations. Moreover, the combination of forecasting development of both in player quality and transfer value gives a comprehensive summary of the most important factors in transfer decisions. In this way, this paper fills a gap in the literature by forecasting the development of model-based indicators for player quality and value on a significant data set in a predictive setting.
@@ 3 Methods
The goal of this study is to find the most suitable machine learning model to forecast the development in player performance with respect to predictive accuracy and uncertainty quantification methods. The assessment of the uncertainty quantification methods is performed using literature on the models and is addressed in the conclusion (section 7). To compare the predictive performance, the models are trained to predict player performance indicators one year ahead in two prediction problems and the corresponding loss values are determined. The two prediction problems concern the prediction of the development in player quality based on the top-down SciSkill rating and the development in monetary value described by the Estimated Transfer Value (ETV) model.
For the player quality prediction problem, data is available from 2014 up to 2022, while the data for the player value covers the years 2016 up to 2021. Because the data of player performance indicators is often dependent over time [26, 16], time-dependent train-test splits are applied to study the predictive performance on unseen data. In both case studies, all data up to 2020 is considered as the training set, and from 2021 and later as the test set. This is visualized in Figure 1. Beforehand, 5% of all data points were left out of both data sets based on stratified sampling for internal studies by the data provider. The root mean square error (RMSE) and mean absolute error (MAE) are determined for these test sets, and performance in the RMSE is considered the most important. The losses are determined on different parts of the test set. First, the loss values of the RMSE and MAE are calculated using the whole test sets. Second, the loss function of the RMSE is also considered for different ages on the test sets because estimating the potential of a player is mostly interesting for young ages. Third, the RMSE is studied on important subgroups of players, like players with large positive or negative development, players with good performance, or players with a high transfer value. By determining the test losses separately on the general population of players, young players, and important subgroups of players, the models are studied on their predictive performance.
@@
3.1 Player quality: SciSkill
The first prediction problem concerns predicting the development in the subsequent year of player quality described by an EM algorithm called the SciSkill [23], which is a generalization of the Elo rating as discussed in section 2. The dataset is restructured to contain monthly data points describing the situation of each player at that time. Each monthly data point consists of the features and the dependent variable. The dependent variable is the difference between the player quality one year ahead and the current player quality value, which is the development in player quality in one year. The features consist of 86 handcrafted features describing for instance the month of year, the current player performance, league strength, time since the most recent game, player characteristics, time series information of the SciSkill, the club’s transfer situation, and the difference in quality between the player and his teammates. The dependent variable and features are described in LABEL:Appendix:_Table_dependent_variable_player_quality and LABEL:Appendix:_Table_features_player_quality.
Only male players with more than 20 games and more than 2 years of data are considered. The final data set consists of 80,568 male professional football players playing in the years 2012 up to 2023. As the data set consists of 3,834,539 data points, there are on average 47.6 monthly data points per player, which corresponds to roughly 4 seasons of data.
3.2 Player value: Estimated Transfer Value
The data set for this prediction problem of the development in player value describes the current situation of professional football players with 58 features. These features consist of indicators of league strength, age, experience, contract situation, the current quality (SciSkill), or the monetary value (ETV) of a football player. The feature set also includes player characteristics, like playing position and age, as well as the differences in quality with teammates, the transfer history of his current club, and league strength. A description of all features and the dependent variable in this prediction problem can be found in LABEL:Appendix:_Table_dependent_variable_player_value and LABEL:Appendix:_Table_features_player_value.
For this prediction problem, the data set consists of biannual data points describing the players’ transfer values in January and July within the period of 2014 up to 2021. Similarly as with the SciSkill, players with less than 20 games, less than 2 years of data, or missing values are excluded. The remaining data set includes 60,175 male professional football players described with 413,177 biannual data points. This corresponds to on average 6.87 data points per player, equivalent to approximately 3.5 seasons of data.
3.3 Models
In this study, linear, tree-based, and kNN-based models are implemented. The linear models are ordinary least squares (OLS), also known as multiple linear regression (MLR), lasso regression, and a linear mixed effect model (LME). For the OLS model [41], feature selection is performed by applying backwards selection with a threshold of the p-values of 0.0001 for the player quality and 0.001 for the player value. These values have been determined via trial and error, and they are smaller than commonly used for significance testing. This is due to the predictive nature of this study and the large data sets combined with the fact that common assumptions of the linear regression model, like normality, do not hold. For lasso regression [42], feature selection is applied by selecting only the variables with nonzero coefficients. The linear mixed effect (LME) model [41] is trained to take into account the influence of a player’s nationality as a random influence. The feature selection for the LME model is done by taking the 20 variables corresponding to the largest absolute value of the coefficients within the lasso model. For these three linear models, no interaction effects are included.
The tree-based models are the decision tree [42], random forest [42], and XGBoost [43] models. For the tree-based models, feature selection is performed by first adding noise variables to the data set and training the models. All features with a larger feature importance than the noise variables are then selected. As tree-based methods can have a bias favoring non-discrete variables, both discrete and continuous variables are added. The discrete features are compared with the discrete noise variables and continuous features with the continuous noise variables.
To investigate the predictive power of time series, three k-nearest neighbors (kNN) models have been implemented to an altered feature set using the Hierarchical Navigable Small Worlds indexer provided by Douze et al. [44]. This feature set consists of time series information with lagged versions of the most important player indicators and is described in LABEL:Appendix:_Table_knn_features_player_quality and LABEL:Appendix:_Table_dependent_variable_player_value in the appendix. First, a normal kNN model is applied to the predictive problem. This model searches for the most similar data points with respect to the Euclidian norm. The prediction is then obtained by applying a weighted average, where closer data points are weighted more heavily as described by Dudani [45]. Possible weighting methods are the reciprocal of the absolute value of the distance, the distance with min-max scaling, or uniform weights. The method of calculating these weights is considered a hyperparameter. The second kNN model is constructed similarly, but it calculates the distances based on the Mahalanobis distance. The Mahalanobis distance projects the features on a decorrelated feature space and calculates the Euclidian distance. This makes it possible to better distinguish differences because the lagged time series features are heavily dependent. Lastly, an adapted RReliefF method [46] is implemented to calculate feature importances of the normal kNN model in a regression context. The features are then multiplied with the feature importances before calculating Euclidian distances to introduce new feature weights. The kNN model is then trained on these reweighted features.
@@ 3.4 Feature importances
The linear and tree-based models have methods that can be used to calculate the feature importances of the models. To investigate what features are important for the development of professional players, the feature importances are calculated. This is similar to the methods by Baouan et al. [39]. However, because the development of the indicator is considered instead of the indicator value itself, it is possible to describe the influence of the indicator value on the development. Min-max scaling is applied to the feature importances to be able to compare them across the different models.
4 Prediction problem on player quality
4.1 General population of players
The loss values on the test set are shown in Figure 4 and Figure 4. The RMSE and MAE agree on the predictive performance of the models to forecast player development in general player quality as they show similar patterns. The results indicate that the XGBoost model attains the lowest loss values, with the random forest having the second-best losses.
@@ 4.1 General population of players
The tree-based models generally obtain the lowest loss values. The difference in performance with the linear models implies that there is some nonlinear or interaction effect in the true underlying relation. On the other hand, the kNN models based on the time series attain the worst loss values. This implies that these models missed out on important information by relying on time series information with a local method. It shows that the development of the player quality is nonlinear and dependent on contextual information.
4.2 Predictions per age
@@
The RMSE loss is also determined for the players of each age in the test set as visualized in Figure 5. The results show that the models predict the development of players between the ages of 23 and 28 most accurately, whereas younger and older players are found to be harder to predict. For the young players, the random forest and XGBoost models provide the most accurate predictions.
The kNN models based on time series perform the worst on the general test set, and Figure 5 shows that this worse performance is especially evident in the RMSE losses corresponding to young players. On the other hand, the tree-based models appear to outperform the linear and kNN-based models for all ages. This explains the better performance on the general test set visible in Figure 4 and Figure 4.
4.3 Prediction on important player groups
The test losses are determined on three different groups of players that are important for the application of the models in this prediction problem: high-quality players, players with a large decrease in performance, and players with a large improvement in performance. The RMSE values on the test set for these subgroups are shown in Figure 6, Figure 8, and Figure 8. In general, the vertical scales show that the RMSE on these subgroups is larger than the one of the general population. This is even more evident for players with a large increase, and it indicates that the development of these players of interest is harder to predict, especially the players with a large increase.
The results also show that the tree-based models are best at predicting the development of all three of these groups of interesting players. The random forest and XGBoost models appear to predict the best for the high-quality players and the players with a large decrease in quality. Figure 8 shows that the XGBoost model outperforms all other models on the group of players with a large increase in quality. The random forest model attains the second-best loss values.
@@ 4.3 Prediction on important player groups
The results indicate that the linear models predict significantly less accurately for the group of players with a large decrease in quality compared to the other models. This shows that interaction effects or nonlinearities are particularly important for predicting a decrease in player quality. On the other hand, the kNN models based on time series predict significantly worse for players with large increases in quality as the large increases are probably for the young players. This is in line with Figure 5, where kNN methods do not perform well on the young players. These differences in losses show that the large increases in player quality are better predicted by global models with contextual information and that decreases are better predicted by methods that include nonlinearities or interaction effects.
4.4 Feature importances
The feature importances of the linear and tree-based models are shown in Figure 9. The feature importances indicate that the age (‘age_years’), age squared (‘age_years_squared’), and the difference between the age and the peak age (‘years_diff_peak_age’) are important features amongst the models. It is known that young players tend to increase in quality, whereas older players often decrease in football quality. Therefore, it is reasonable that these age-related factors are indeed important in the development of the quality of professional football players.
Multiple models indicate that the player quality (‘sciskill’) at the time of prediction, the difference between the player quality and the average quality in the team (‘sciskill_diff_mean_team’), and the number of months since the last registered game (‘previous_zero_months’) are important features. The importance of the player quality indicates that players from different qualities show different development patterns. The influence of the difference in quality between the player and the team is caused by players having good performances when their teams play well. Consequently, the player’s quality grows towards the average team quality, which explains the importance of the difference in quality between the player and his team. Next to that, the SciSkill model penalizes players when they have not played games for a long time. As this penalty is applied after the next game of a player, the quality of a player can be expected to decrease when the number of months since the last game is large. The importance of the number of months since the last registered game shows that the models can find this pattern. In this way, the feature importance can be used to identify what factors are important for the development of professional football players and to test whether models behave as expected.
@@ 4.4 Feature importances
5 Prediction problem on player value
5.1 General population of players

### v2

@@ 2.3 Predicting future values
Although the studies discussed above are concerned with the quantification of the quality and monetary value of football players at the present moment, only limited work has been done on the future development of player performance. Apostolou and Tjortjis [42] have conducted a small-scale study with the aim of predicting the number of future goals of the two football players Lionel Messi and Luis Suárez using a random forest, logistic regression, a multi-layer perception classifier, and a linear support vector classifier. Pantzalis and Tjortjis [43] have predicted the expert ratings in the next season of 59 center-backs in the English Premier League based on one season of player attributes from a popular football manager simulation game. Their method uses a linear regression model to describe the in-sample patterns. Giannakoulas et al. [44] have trained linear regression, random forest, and multi-layer perceptron models to predict the number of goals of a football player in a season before the start of the corresponding season. Their dataset entails around 800 football players. Similarly, the models by Markopoulou et al. [45] have been trained to predict the number of goals by looking at the creation of different models per competition. This study on 424 football players shows that the best results are often obtained using XGBoost models for this prediction problem and that making different models for different competitions might be beneficial.
@@
Our research treats forecasting of both player quality and transfer value. The current paper builds on the existing literature about forecast player performance by studying the long-term forecasting of a model-based player quality indicator on a larger data set. Additionally, this research avoids the bias in current literature that better describes offensive players by using a more general top-down rating. Our research contributes to the existing knowledge of forecasting the monetary value because the models are trained to perform out-of-sample prediction, which makes it possible to apply them to unseen situations. Moreover, the combination of forecasting the development of both in player quality and transfer value gives a comprehensive summary of the most important factors in transfer decisions. In this way, this paper fills a gap in the literature by forecasting the development of model-based indicators for player quality and value on a significant data set in a predictive setting.
@@ 3 Methods
The goal of this study is to find the most suitable machine learning model to forecast the development in player performance with respect to predictive accuracy and uncertainty quantification methods. The assessment of the uncertainty quantification methods was performed using the literature on the models in section 2. To compare the predictive performance, the models are trained to predict player performance indicators one year ahead in two prediction problems, and the corresponding loss values are determined. The two prediction problems concern the prediction of the development in player quality based on the top-down SciSkill rating and the development in monetary value described by the Estimated Transfer Value (ETV) model.
3.1 Data
Two data sets were obtained to study the predictive performance of the models in this paper. The features of the data points in both prediction problems represent the historical situation of professional football players. The corresponding label is the player performance indicator that was recorded one year later. The data availability for different years is visualized in Figure 2. For the player quality prediction problem, data is available from 2014 up to 2022, while the data for the player value covers the years 2016 up to 2021.
@@ 3.1 Data
3.1.1 Player quality: SciSkill
The first prediction problem concerns predicting the development in the subsequent year of the player quality described by an EM algorithm called the SciSkill [23], which is a generalization of the Elo rating as discussed in section 2. The dataset is restructured to contain monthly data points describing the situation of each player at that time. Each monthly data point consists of the features and the dependent variable. The dependent variable is the difference between the player quality one year ahead and the current player quality value, which is the development in player quality in one year. The features set consist of 86 features constructed with domain knowledge describing for instance the month of year, the current player performance, league strength, time since the most recent game, player characteristics, time series information of the SciSkill, the club’s transfer situation, and the difference in quality between the player and his teammates. The dependent variable and features are described in LABEL:Appendix:_Table_dependent_variable_player_quality and LABEL:Appendix:_Table_features_player_quality.
Only male players with more than 20 games and more than 2 years of data are considered. The final data set consists of 80,568 male professional football players playing in the years 2012 to 2023. As the data set consists of 3,834,539 data points, there are on average 47.6 monthly data points per player, which corresponds to roughly 4 seasons of data.
3.1.2 Player value: Estimated Transfer Value
The data set for this prediction problem of the development in player value describes the current situation of professional football players with 58 features. These features were constructed based on domain knowledge and data availability. Because there were fewer data points than for the player quality case study, the number of features was reduced by eliminating highly correlated features. These features consist of indicators of league strength, age, experience, contract situation, the current quality (SciSkill), or the monetary value (ETV) of a football player. The feature set also includes player characteristics, like playing position and age, as well as the differences in quality with teammates, the transfer history of his current club, and league strength. A description of all features and the dependent variable in this prediction problem can be found in LABEL:Appendix:_Table_dependent_variable_player_value and LABEL:Appendix:_Table_features_player_value.
For this prediction problem, the data set consists of biannual data points describing the players’ transfer values in January and July within the period from 2014 to 2021. Similarly to the SciSkill, players with less than 20 games, less than 2 years of data, or missing values are excluded. The remaining data set includes 60,175 male professional football players described with 413,177 biannual data points. This corresponds to an average of 6.87 data points per player, equivalent to approximately 3.5 seasons of data.
3.2 Model implementation
3.2.1 Model assessment
Supervised models are trained on the prediction problems of player quality and player value. The root mean square error (RMSE) and mean absolute error (MAE) are determined for the test set. The RMSE is more strongly influenced by large errors, which can be expected to happen for ‘superstar’ players. Since these players are especially of interest to practitioners, the performance in the RMSE is considered the most important. The losses are determined on different parts of the test set. First, the loss values of the RMSE and MAE are calculated using the complete test sets. Second, the loss function of the RMSE is also considered for different age groups on the test sets because estimating the potential of a player is mostly interesting for young players. Third, the RMSE is studied on important subgroups of players, like players with large positive or negative development, players with good performance, or players with a high transfer value. By determining the test losses separately on the general population of players, young players, and important subgroups of players, the models are studied on their predictive performance.
Beforehand, 5% of all football players are left out of both data sets based on stratified sampling for internal studies by the data provider. Because the data of player performance indicators is often dependent on time [26, 16], time-dependent train-test splits are applied to study the predictive performance on unseen data as visualized in Figure 2. In both case studies, all data up to 2020 is considered as the training set, and from 2021 and later as the test set.
3.2.2 Model training
In this study, linear, tree-based, and kNN-based models are implemented. The linear models are ordinary least squares (OLS), also known as multiple linear regression (MLR), lasso regression, and a linear mixed effect model (LME). For the OLS model [48], feature selection is performed by applying backwards selection with a threshold of the p-values of 0.0001 for the player quality and 0.001 for the player value. These values have been determined via trial and error, considering values of $10^{k}$ for multiple values of $k$, and they are smaller than commonly used for significance testing. This is due to the predictive nature of this study and the large data sets combined with the fact that common assumptions of the linear regression model, like normality, do not hold. For lasso regression [49], feature selection is applied by selecting only the variables with nonzero coefficients. The linear mixed effect (LME) model [48] is trained to take into account the influence of a player’s nationality as a random influence. The feature selection for the LME model is done by taking the 20 variables corresponding to the largest feature importances within the lasso model. This is done for computational feasibility. For these three linear models, no interaction effects are included.
The tree-based models are the decision tree [49], random forest [49], and XGBoost [30] models. For the tree-based models, feature selection is performed by first adding noise variables to the data set and training the models. All features with a larger feature importance than the noise variables are then selected. As tree-based methods can have a bias favoring non-discrete variables, both discrete and continuous variables are added. The discrete features are compared with the discrete noise variables, and the continuous features with the continuous noise variables.
To investigate the predictive power of time series, three k-nearest neighbors (kNN) models have been implemented using the Hierarchical Navigable Small Worlds indexer provided by Douze et al. [50]. This was done on an altered feature set consisting of time series information with lagged versions of the most important player indicators. Because the kNN model suffers from the curse of dimensionality [27, Section 2.5], the feature set is kept relatively small. The features and labels of this altered data set are described in LABEL:Appendix:_Table_knn_features_player_quality and LABEL:Appendix:_Table_dependent_variable_player_value in the appendix. First, a normal kNN model is applied to the predictive problem. This model searches for the most similar data points with respect to the Euclidean norm. The prediction is then obtained by applying a weighted average, where closer data points are weighted more heavily as described by Dudani [31]. Possible weighting methods are the reciprocal of the absolute value of the distance, the distance with min-max scaling, or uniform weights. The method of calculating these weights is considered a hyperparameter. The second kNN model is constructed similarly, but it calculates the distances based on the Mahalanobis distance. The Mahalanobis distance projects the features on a decorrelated feature space and calculates the Euclidean distance. This makes it possible to better distinguish differences because the lagged time series features are heavily dependent. Lastly, an adapted RReliefF method [51] is implemented to calculate feature importances of the normal kNN model in a regression context. The features are then multiplied by the feature importances before calculating Euclidean distances to introduce new feature weights. The kNN model is then trained on these reweighted features.
3.2.3 Hyperparameter tuning
@@ 3.3 Feature importances
To illustrate the advantages of using explainable models, the feature importances are studied in both predictive problems. Among the models in this paper, the linear and tree-based models have methods that can be used to calculate the feature importances of the models. To investigate what features are important for the development of professional players, the feature importances are calculated. This is similar to the methods by Baouan et al. [47]. However, because the development of the indicator is considered instead of the indicator value itself, it is possible to describe the influence of the indicator value on the development. Min-max scaling is applied to the feature importances to be able to compare them across the different models.
4 Results
4.1 Prediction problem on player quality
4.1.1 General population of players
The loss values on the test set are shown in Figure 5 and Figure 5. The RMSE and MAE agree on the predictive performance of the models to forecast player development in general player quality, as they show similar patterns. The results indicate that the XGBoost model attains the lowest loss values, with the random forest having the second-best losses.
@@ 4.1.1 General population of players
The tree-based models generally obtain the lowest loss values. The difference in performance with the linear models implies that there is some nonlinear or interaction effect in the true underlying relation. On the other hand, the kNN models based on the time series attain the worst loss values. This implies that these kNN-based models missed out on important information by relying on time series information with a local method. It shows that the development of player quality is nonlinear and dependent on contextual information.
4.1.2 Predictions per age
@@
The RMSE loss is also determined for the players of each age in the test set as visualized in Figure 6. The XGBoost and random forest models perform better for all ages. For ages above 25, the XGBoost model has better loss values compared to the random forest model, whereas the differences are less evident for the lower ages. Because the younger ages are most important for estimating the potential of players in practice, these results indicate that both the XGBoost and random forest models are favorable models.
In general, Figure 6 shows that all models predict best on players with ages 24 up to 28. The quality of younger or older players is more volatile because young players tend to increase in quality, and older players tend to decrease in quality. Additionally, more data is available on players around their peak age. The worse predictive performance on the younger and older players is explained by the lower number of data points combined with a higher volatility in player quality.
Additionally, the results show that the kNN models perform worse on young players. The linear methods tend to perform worse for older players. This indicates that nonlocal methods with more features beyond time series perform better for young ages, whereas the patterns for older players involve nonlinearities or interaction effects. This explains the low loss values of the tree-based models on the general test set as shown in Figure 5 and Figure 5.
4.1.3 Prediction on important player groups
The test losses are determined on three different groups of players that are important for the application of the models in this prediction problem: high-quality players, players with a large decrease in performance, and players with a large improvement in performance. The RMSE values on the test set for these subgroups are shown in Figure 7, Figure 9, and Figure 9. In general, the vertical scales show that the RMSE on these subgroups is larger than that of the general population. This is even more evident for players with a large increase, and it indicates that the development of these players of interest is harder to predict, especially the players with a large increase.
The results also show that the tree-based models are best at predicting the development of all three of these subgroups of players. The random forest and XGBoost models appear to predict the best for the high-quality players and the players with a large decrease in quality. Figure 9 shows that the XGBoost model outperforms all other models on the group of players with a large increase in quality. The random forest model attains the second-best loss values.
@@ 4.1.3 Prediction on important player groups
The results indicate that the linear models predict significantly less accurately for the group of players with a large decrease in quality compared to the other models. This shows that interaction effects or nonlinearities are particularly important for predicting a decrease in player quality. On the other hand, the kNN models based on time series predict significantly worse for players with large increases in quality, as the large increases are probably for the young players. This is in line with Figure 6, where kNN methods do not perform well on the young players. These differences in losses show that the large increases in player quality are better predicted by global models with contextual information and that decreases are better predicted by methods that include nonlinearities or interaction effects.
4.1.4 Feature importances
The feature importances of the linear and tree-based models are shown in Figure 10. The feature importances indicate that the age (‘age_years’), age squared (‘age_years_squared’), and the difference between the age and the peak age (‘years_diff_peak_age’) are important features amongst the models. It is known that young players tend to improve in quality, whereas older players often decrease in quality. Therefore, it is reasonable that these age-related factors are indeed important in the development of the quality of professional football players.
@@ 4.1.4 Feature importances
Next to that, other features have also been assigned a large importance by the models. For instance, the player quality at prediction time (‘sciskill’) has a large feature importance. This indicates that the development patterns depend on the player level, which is reflected by the larger loss values in Figure 7 compared to Figure 5. The importance of the difference between player quality and average team quality (‘sciskill_diff_mean_team’) can be explained by the fact that players tend to have good performances when their team plays well. Consequently, the player’s quality grows towards the average team quality, which explains the importance of the difference in quality between the player and his team. Lastly, the feature importance assigns a large importance to the number of months since the last registered game (‘previous_zero_months’). The SciSkill model penalizes players when they have not played games for a long time. As this penalty is applied after the next game of a player, the quality of a player can be expected to decrease when the number of months since the last game is large. The importance of the number of months since the last registered game shows that the models can find this pattern.
To summarize, the models indicate that the player age, player quality, the difference in quality between player and team, and the number of months since the last game are the most important features to predict future player quality. In this way, the feature importance can be used to identify what factors are important for the development of professional football players and to test whether models behave as expected.
4.1.5 Predictive accuracy
The results of training the models to predict the future player quality show that the XGBoost model provides the most accurate predictions on the general population of players. The XGBoost model outperforms the other models on young players, high-quality players, and players with either a large decrease or increase in football quality. The random forest model predicts second best, with similar performances on the young players and most of these interesting subgroups of players. For the problem of predicting a football player’s development in quality, it is concluded that the best predictive performance is attained by the XGBoost model, followed closely by the random forest model.
4.2 Prediction problem on player value
4.2.1 General population of players